Assume that X and Y are discrete random variables having the joint pmf given by the following chart

                             Y

               0            1             2

       0     0.1         0.1          0.3

X    1     0.3         0.1          0.1

              

a. Find the probability that Y is greater than X.

b. Find the covariance between X and Y.
Answer #1

Toy 0.3 । ory 0.2 ony 1 P ( xo y XV) = P(x=0, y = 4). Selon +P/x= 12 Y=2) 0.1 + 0.1 50.2, O EXT=x 0:3)+{1x05) - 054 hay until